OpenStudy (anonymous):

Find the exact value by using a half-angle identity. cos(-pi/8)

OpenStudy (anonymous):

\[\cos (-\frac{ \pi }{ 8 })\]

OpenStudy (zehanz):

Do you know a formula for the half-angle?

OpenStudy (anonymous):

I don't have it written down with my others...

OpenStudy (zehanz):

Maybe you are familiar with: \(\cos(2\theta)=2\cos^{2}(\theta)-1 \)

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ \cfrac{1}{8}\cdot 2\implies \cfrac{1}{4}}} \\ \quad \\ \quad \\ cos\left(-\frac{\pi}{8}\right)\implies cos\left(\cfrac{{\color{blue}{ \frac{-\pi}{4}}}}{2}\right)=\pm\sqrt{\cfrac{1+cos\left({\color{blue}{ \frac{\pi}{4}}}\right)}{2}}\)

OpenStudy (jdoe0001):

hmm \(\bf {\color{blue}{ \cfrac{1}{8}\cdot 2\implies \cfrac{1}{4}}} \\ \quad \\ \quad \\ cos\left(-\frac{\pi}{8}\right)\implies cos\left(\cfrac{{\color{blue}{ \frac{-\pi}{4}}}}{2}\right)=\pm\sqrt{\cfrac{1+cos\left({-\color{blue}{ \frac{\pi}{4}}}\right)}{2}}\)

OpenStudy (jdoe0001):

or \(\bf {\color{blue}{ \cfrac{1}{8}\cdot 2\implies \cfrac{1}{4}}} \\ \quad \\ \quad \\ cos\left(-\frac{\pi}{8}\right)\implies cos\left(\cfrac{{\color{blue}{ \frac{-\pi}{4}}}}{2}\right)=\pm\sqrt{\cfrac{1+cos\left({\color{blue}{ \frac{-\pi}{4}}}\right)}{2}}\) anyhow, that's a well known angle that'd be in your Unit Circle

OpenStudy (anonymous):

None of my choices look like that.. :( \[A) \frac{ 1 }{ 2 } \sqrt{2-\sqrt{2}}\] \[B) \frac{ 1 }{ 2 } \sqrt{1-\sqrt{2}}\] \[C) \frac{ 1 }{ 2 } \sqrt{2 + \sqrt{2}}\] \[D) \frac{ 1 }{ 2 } \sqrt{1+\sqrt{2}}\]

myininaya (myininaya):

you are suppose to simplify what joe what

OpenStudy (anonymous):

Do what?

OpenStudy (anonymous):

@jdoe0001

myininaya (myininaya):

I mean wrote

myininaya (myininaya):

you are suppose to simplify what he wrote

myininaya (myininaya):

what is cos(-pi/4) used the fact that it is on the unit circle to find the what cos(-pi/4) is

OpenStudy (jdoe0001):

\( \bf {\color{blue}{ \cfrac{1}{8}\cdot 2\implies \cfrac{1}{4}}} \\ \quad \\ \quad \\ cos\left(-\frac{\pi}{8}\right)\implies cos\left(\cfrac{{\color{blue}{ \frac{-\pi}{4}}}}{2}\right)=\pm\sqrt{\cfrac{1+cos\left({\color{blue}{ \frac{-\pi}{4}}}\right)}{2}} \\ \quad \\ \textit{so what's the }cos\left({\color{blue}{ \frac{-\pi}{4}}}\right)\quad ?\)

OpenStudy (anonymous):

-sqrt2/2?

OpenStudy (jdoe0001):



OpenStudy (anonymous):

So it's positive?

OpenStudy (jdoe0001):

anyhow, yes, but is a POSITIVE amount, since in the IV Quadrant the cosine is positive, thus \(\bf cos\left(-\frac{\pi}{8}\right)\implies cos\left(\cfrac{{\color{blue}{ \frac{-\pi}{4}}}}{2}\right)=\pm\sqrt{\cfrac{1+cos\left({\color{blue}{ \frac{-\pi}{4}}}\right)}{2}} \\ \quad \\ \implies cos\left(\cfrac{{\color{blue}{ \frac{-\pi}{4}}}}{2}\right)=\pm\sqrt{\cfrac{1+\frac{\sqrt{2}}{2}}{2}}\)

OpenStudy (jdoe0001):

then just simplify the rational

OpenStudy (anonymous):

What does that mean?

OpenStudy (jdoe0001):

h..... does . what mean? simplify?

myininaya (myininaya):

Teachers don't normally allow compound fractions in answers. When we say simplify one of the things you need to make sure of is that there are no compound fractions in your answer.

OpenStudy (jdoe0001):

right, make it a simple fraction inside the radical

OpenStudy (anonymous):

I don't get how the 1/2 gets to the outside...